OpenStudy (anonymous):

Solve for x: 3/4x + 5/8 = 4x

OpenStudy (math&ing001):

First get all the x-terms in one side like this : 5/8 = 4x - 3/4x Then factor by x : 5/8 = (4 - 3/4)x Then divide by (4 - 3/4).

OpenStudy (anonymous):

Yes I was going to say that

OpenStudy (solomonzelman):

in 3/4x is x part of the denominator or not ?